OpenStudy (anonymous):

A 1.5-kg cart is pulled with a force of 7.3 N at an angle of 40 degrees above the horizontal. If a kinetic friction force of 3.2 N acts against the motion, the cart's acceleration along the horizontal surface will be?

OpenStudy (anonymous):

This is a simple balance of forces.\[\sum F = m*a = F - F_f\]

OpenStudy (anonymous):

what is the F=? and Ff=?

OpenStudy (anonymous):

i don't get it..

OpenStudy (anonymous):

Okay. From Newton's first law, The net force equals the mass times the acceleration. F is the applied force, F_f is the frictional force. Since frictional force always opposes the motion, it must be directed in the opposite direction as the applied force.

OpenStudy (anonymous):

Newton's second law rather.

OpenStudy (anonymous):

yes. Oh.. so it's Fa - Ff = m*a = F

OpenStudy (anonymous):

what's the angel's role?

OpenStudy (anonymous):

It is just\[m*a = F_a - F_f\]

OpenStudy (anonymous):

ohh.

OpenStudy (anonymous):

Ah. I misread the problem. We need to find the component of the applied force that is directed along the path of motion. \[F_p = F_a \sin(\theta)\]\[m*a = F_p - F_f\]

OpenStudy (anonymous):

what is Fp?

OpenStudy (anonymous):

I told you what F_p is. \[F_p = F_a \sin(\theta)\]It is the component of the force that acts along the path of motion.

OpenStudy (anonymous):

oh but like is it like applied force..

OpenStudy (anonymous):

Fp = 7.3 N sin(40 = 5.44 m*a = Fp - Ff 1.5*a = 5.44 - 3.2 1.5*a = 2.24 a = 1.49?

OpenStudy (anonymous):

Yes.

OpenStudy (anonymous):

Do you understand why?

OpenStudy (anonymous):

no, I don't. Plus it's not in the answer provided.. please explain D: I'm so bad at physics i don't even know why i took it >_<"

OpenStudy (anonymous):

First, begin every problem like this with a free body diagram. Now, we need to break up the applied force into its normal components. That is, x and y based on the drawn coordinate system (a coordinate system should be included in every free body diagram)\[F_x = F_a \cos(\theta)\]\[F_y = F_a \sin(\theta)\](Whoops! I gave you the wrong expression for the applied force. I'm sorry.) We are interested in the horizontal acceleration. Or the acceleration in the x-direction. Now, let's write a balance of forces. \[m*a_x = F_a \cos(\theta) - F_f\]